Circular No. 5947

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) For the Hubble Space Telescope observing team, H. A. Weaver, Space

Telescope Science Institute, reports the following results from their analysis of HST images taken during Jan. 24–27: “The HST data can only place upper limits on the sizes of the cometary fragments, since there is no clear evidence of a point source near any of the condensations. The coma spatial brightness distribution cannot be fitted by a simple analytical (or numerical) model, so there is no reliable way of judging how much of the brightness is due to an unresolved source. If one is willing to accept that the coma becomes completely flat at the position of a fragment, then both the July 1993 (Weaver et al. 1994, Science 263, 787) and Jan. 1994 HST data indicate that the largest fragment could be as large as 4 km in diameter. However, since it seems unlikely for the coma profile to flatten out like this, and since the new HST data show that the coma between 0′′.1 and 0′′.6 is reasonably well approximated by a power law with an index of −1.33, it is possible that all of the light observed (to within the noise level) is due to the coma. In the new images, most of the condensations have a coma that is circularly symmetric out to ∼ 0′′.6 from the peak pixel; then there is a strong divergence in the brightness distribution on one side of the ‘train’ versus the other. This behavior is also present in the July 1993 HST data. The spatial brightness profiles have become significantly steeper in the new data compared to the old (after taking into account the spherical aberration present in the old data), indicating that there has been significant change in the dust distribution between July 1993 and Jan. 1994. There have also been some changes in both the relative and absolute magnitudes of the components. Fragment 15 is now ∼ 15 percent brighter than fragment 7a. Fragments 5, 8a and 8b have markedly different spatial brightness distributions from most of the others (but not all condensations have been examined in detail). There is a brightness spur on fragment 5 that extends into the ‘dust-free’ side of the train and that may indicate that this nucleus is fragmenting further. Nine of the fragments are significantly displaced from the train, with all displacements being into the dusty side. The region around fragment 8a looks more like a ‘streak’ than a strong condensation, possibly indicating that there is only a swarm of material in this region, rather than dust plus a dominant fragment. The case of 8b seems to be intermediate between that of 8a and the more well-condensed cases.”

1994 March 10 Brian G. Marsden